OpenStudy (anonymous):

WILL GIVE FANS AND METALS!!! -write an equation in point-slope form of the line through point J(4,1)with slope -4.

OpenStudy (anonymous):

The formula for point-slope form is: y – y1 = m(x – x1) So, all we need to do is plug in the points and slope. (4, 1) 4 is x1. 1 is y1. And of course, -4 is m. Lets plug it in. y - 1 = -4(x - 4) <----- Answer. (:

OpenStudy (anonymous):

can u help me with more? @BunnyBree

OpenStudy (anonymous):

I'll try.

OpenStudy (anonymous):

lines a,b,c,and d are in the in the same plane. Line a is parallel to line c. line b and c are perpendicular to line d. Based on this,tell how lines a and b are related. justify your answer. @BunnyBree

OpenStudy (anonymous):

@BunnyBree hello?

OpenStudy (anonymous):

Sorry, I was busy doing something for school. c: A ∥ B because B and C are both ⊥ to the same line, D, they are ∥ to each other. Because A ∥ C, by the Transitive Property of Parallel Lines, A ∥ B.